# MedVAE: Efficient Automated Interpretation of Medical Images with Large-Scale Generalizable Autoencoders

### v1

@@ 3   Discussion
Our study presents several opportunities for future work. First, additional research into model architectures, data augmentation architectures, and training strategies would be useful for building effective downstream CAD models that can learn from latent representations. In addition, the batch size and efficiency benefits afforded by latent representations raise the possibility of training large-scale foundation models using downsized latent representations. Whereas foundation models traditionally require significant computational resources and training time, utilizing downsized latent representations that preserve diagnostic features can greatly accelerate model training, particularly in resource-constrained settings. Future work can explore foundation model performance and scaling laws in this context. Finally, future work can explore additional autoencoder training strategies to better preserve clinically-relevant features at high downsizing factors.
@@ 4.3 Training autoencoders for medical images
In the context of 2D imaging modalities, the second training stage takes the form of a lightweight fine-tuning procedure designed to maximize consistency in clinically-relevant features between the input image and the latent representation. Our key insight here is that image embeddings generated by BiomedCLIP [55] can effectively capture clinically-relevant features in 2D medical images, suggesting utility as a guidance mechanism during training. We freeze all parameters in the encoder and decoder of the VAE. During training, the input image $x_{i}$ is passed through the frozen VAE encoder to generate the latent representation $z_{i}$; then, $z_{i}$ is passed through a series of lightweight, trainable projection layers, which yield an output representation $\bar{z_{i}}$ with the same size as $z_{i}$. Let the function $b(\cdot)$ represent the BiomedCLIP embedding function. We optimize the projection layer weights using a domain-specific embedding consistency loss, which takes the form of an $L_{2}$ loss between $b(x_{i})$ and $b(\bar{z_{i}})$. All downstream evaluations of latent representation quality are performed with the projected latent $\bar{z_{i}}$. We perform Stage 2 training using the curated 2D training dataset with one million images. Our procedure yields four 2D MedVAE autoencoders with various downsizing factors and number of latent channels:
@@ Acknowledgments
MV is supported by graduate fellowship awards from the Department of Defense (NDSEG), the Knight-Hennessy Scholars program at Stanford University, and the Quad program. AK is supported by graduate fellowships from Tau Beta Pi and the Knight-Hennessy Scholars program at Stanford University. RS was supported by the Rubicon fellowship of the Dutch National Research Council (NWO). This work was supported in part by NIH grants R01 HL155410, R01 HL157235, by AHRQ grant R18HS026886, and by the Gordon and Betty Moore Foundation. CL is supported by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through The Advanced Research Projects Agency for Health (ARPA-H). AC is supported by NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060, AY2AX000045, and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021.

### v2

@@ 2.4 Reconstructed image quality
2.5 Ablations
We further ablate the inclusion of the embedding consistency loss term in the Stage 1 training procedure. We find that the embedding consistency loss term helps improve reconstructed image quality, particularly at lower compression factors. For instance, at a compression factor of $f=16$, Stage 1 training without the embedding consistency loss term achieves a PSNR of $37.27_{\pm 0.08}$ and an MS-SSIM of $0.992_{\pm 0.0}$ on mammograms. In comparison, Stage 1 training with the embedding consistency loss term achieves a PSNR of $37.57_{\pm 0.08}$ and an MS-SSIM of $0.993_{\pm 0.0}$, as shown in Table 3.
@@ 3   Discussion
Our study presents several opportunities for future work. First, additional research into model architectures, data augmentation approaches, and training strategies would be useful for building effective downstream CAD models that can learn from latent representations. In addition, the batch size and efficiency benefits afforded by latent representations raise the possibility of training large-scale foundation models using downsized latent representations. Whereas foundation models traditionally require significant computational resources and training time, utilizing downsized latent representations that preserve diagnostic features can greatly accelerate model training, particularly in resource-constrained settings. Future work can explore foundation model performance and scaling laws in this context. Finally, future work can explore additional autoencoder training strategies to better preserve clinically-relevant features at high downsizing factors.
@@ 4.3 Training autoencoders for medical images
In the context of 2D imaging modalities, the second training stage takes the form of a lightweight fine-tuning procedure designed to maximize consistency in clinically-relevant features between the input image and the latent representation. Our key insight here is that image embeddings generated by BiomedCLIP [55] can effectively capture clinically-relevant features in 2D medical images, suggesting utility as a guidance mechanism during training [^1] We use the BiomedCLIP-PubMedBERT_256-vit_base_patch16_224 model available on HuggingFace at https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224.. We freeze all parameters in the encoder and decoder of the VAE. During training, the input image $x_{i}$ is passed through the frozen VAE encoder to generate the latent representation $z_{i}$; then, $z_{i}$ is passed through a series of lightweight, trainable projection layers, which yield an output representation $\bar{z_{i}}$ with the same size as $z_{i}$. Let the function $b(\cdot)$ represent the BiomedCLIP embedding function. We optimize the projection layer weights using a domain-specific embedding consistency loss, which takes the form of an $L_{2}$ loss between $b(x_{i})$ and $b(\bar{z_{i}})$. All downstream evaluations of latent representation quality are performed with the projected latent $\bar{z_{i}}$. We perform Stage 2 training using the curated 2D training dataset with one million images. Our procedure yields four 2D MedVAE autoencoders with various downsizing factors and number of latent channels:
@@ Acknowledgments
MV is supported by graduate fellowship awards from the Department of Defense (NDSEG), the Knight-Hennessy Scholars program at Stanford University, and the Quad program. AK is supported by graduate fellowships from the Knight-Hennessy Scholars program at Stanford University and the Tau Beta Pi society. RS was supported by the Rubicon fellowship of the Dutch National Research Council (NWO). AC is supported by NIH grants R01 HL167974, R01HL169345, R01 AR077604, R01 EB002524, R01 AR079431, P41 EB027060, AY2 AX000045, and 1AYS AX0000024-01; ARPA-H grants AY2AX000045 and 1AYSAX0000024-01; and NIH contracts 75N92020C00008 and 75N92020C00021. A.C. has provided consulting services to Patient Square Capital, Chondrometrics GmbH, and Elucid Bioimaging; is co-founder of Cognita; has equity interest in Cognita, Subtle Medical, LVIS Corp, Brain Key. CL is supported by NIH grants R01 HL155410, R01 HL157235, by AHRQ grant R18HS026886, and by the Gordon and Betty Moore Foundation. CL is also supported by the Medical Imaging and Data Resource Center (MIDRC), which is funded by the National Institute of Biomedical Imaging and Bioengineering (NIBIB) under contract 75N92020C00021 and through the Advanced Research Projects Agency for Health (ARPA-H).
This research was funded, in part, by the Advanced Research Projects Agency for Health (ARPA-H). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.